# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ Appendix F Algorithms
mix box_count_threshold thresholded lac additive 23.9744 0.0272 2.9348 0.0940 12.9752 0.0968 0.1728
mix box_count_threshold thresholded lac multiplicative 23.9744 0.0272 1.8538 0.1016 12.9752 0.0968 0.1830
mix box_count_threshold thresholded aps additive 23.9744 0.0272 2.9348 0.0924 10.6293 0.0945 0.1618
mix box_count_threshold thresholded aps multiplicative 23.9744 0.0272 1.6141 0.0984 10.2883 0.0953 0.1797
mix box_count_threshold pixelwise lac additive 23.9744 0.0272 1.0432 0.0984 0.9060 0.1000 0.1736
mix box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 0.9998 0.1016 0.9060 0.1000 0.1830
mix box_count_threshold pixelwise aps additive 23.9744 0.0272 1.0432 0.0984 0.9937 0.0724 0.1500
mix box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 0.9998 0.1016 0.9772 0.0724 0.1541
mix box_count_threshold boxwise lac additive 23.9744 0.0272 1.3962 0.0984 0.9060 0.1000 0.1736
mix box_count_threshold boxwise lac multiplicative 23.9744 0.0272 1.0432 0.1016 0.9060 0.1000 0.1830
mix box_count_threshold boxwise aps additive 23.9744 0.0272 1.3962 0.0984 0.9937 0.0724 0.1500
mix box_count_threshold boxwise aps multiplicative 23.9744 0.0272 1.0432 0.1016 0.9772 0.0724 0.1541
mix box_count_recall thresholded lac additive 14.4056 0.0262 2.9348 0.0940 12.9752 0.0968 0.1728
mix box_count_recall thresholded lac multiplicative 14.4056 0.0262 1.8538 0.1016 12.9752 0.0968 0.1830
mix box_count_recall thresholded aps additive 14.4056 0.0262 2.9348 0.0924 10.6293 0.0945 0.1618
mix box_count_recall thresholded aps multiplicative 14.4056 0.0262 1.6141 0.0984 10.2883 0.0953 0.1797
mix box_count_recall pixelwise lac additive 14.4056 0.0262 1.0432 0.0984 0.9060 0.1000 0.1736
mix box_count_recall pixelwise lac multiplicative 14.4056 0.0262 0.9998 0.1016 0.9060 0.1000 0.1830
mix box_count_recall pixelwise aps additive 14.4056 0.0262 1.0432 0.0984 0.9937 0.0724 0.1500
mix box_count_recall pixelwise aps multiplicative 14.4056 0.0262 0.9998 0.1016 0.9772 0.0724 0.1541
mix box_count_recall boxwise lac additive 14.4056 0.0262 1.3962 0.0984 0.9060 0.1000 0.1736
mix box_count_recall boxwise lac multiplicative 14.4056 0.0262 1.0432 0.1016 0.9060 0.1000 0.1830
mix box_count_recall boxwise aps additive 14.4056 0.0262 1.3962 0.0984 0.9937 0.0724 0.1500
mix box_count_recall boxwise aps multiplicative 14.4056 0.0262 1.0432 0.1016 0.9772 0.0724 0.1541
giou box_count_threshold thresholded lac additive 23.9744 0.0272 28.1254 0.0932 36.1885 0.0976 0.1708
giou box_count_threshold thresholded lac multiplicative 23.9744 0.0272 24.4546 0.1000 36.1885 0.0976 0.1830
giou box_count_threshold thresholded aps additive 23.9744 0.0272 28.1254 0.0924 34.8062 0.0953 0.1602
giou box_count_threshold thresholded aps multiplicative 23.9744 0.0272 23.8266 0.0984 35.1776 0.0968 0.1805
giou box_count_threshold pixelwise lac additive 23.9744 0.0272 25.1563 0.0952 35.0695 0.1000 0.1716
giou box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 24.2717 0.1000 35.0695 0.1000 0.1830
giou box_count_threshold pixelwise aps additive 23.9744 0.0272 25.1563 0.0952 34.8062 0.0953 0.1602
giou box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 23.9725 0.0984 35.1776 0.0968 0.1805
giou box_count_threshold boxwise lac additive 23.9744 0.0272 28.6734 0.0952 35.4345 0.0992 0.1716
giou box_count_threshold boxwise lac multiplicative 23.9744 0.0272 25.0144 0.1000 35.4345 0.0992 0.1830
giou box_count_threshold boxwise aps additive 23.9744 0.0272 28.6734 0.0952 34.8062 0.0953 0.1602
giou box_count_threshold boxwise aps multiplicative 23.9744 0.0272 24.8396 0.0984 35.1776 0.0968 0.1805
giou box_count_recall thresholded lac additive 14.4056 0.0262 28.1254 0.0932 36.1885 0.0976 0.1708
giou box_count_recall thresholded lac multiplicative 14.4056 0.0262 24.4546 0.1000 36.1885 0.0976 0.1830
giou box_count_recall thresholded aps additive 14.4056 0.0262 28.1254 0.0924 34.8062 0.0953 0.1602
giou box_count_recall thresholded aps multiplicative 14.4056 0.0262 23.8266 0.0984 35.1776 0.0968 0.1805
giou box_count_recall pixelwise lac additive 14.4056 0.0262 25.1563 0.0952 35.0695 0.1000 0.1716
giou box_count_recall pixelwise lac multiplicative 14.4056 0.0262 24.2717 0.1000 35.0695 0.1000 0.1830
giou box_count_recall pixelwise aps additive 14.4056 0.0262 25.1563 0.0952 34.8062 0.0953 0.1602
giou box_count_recall pixelwise aps multiplicative 14.4056 0.0262 23.9725 0.0984 35.1776 0.0968 0.1805
giou box_count_recall boxwise lac additive 14.4056 0.0262 28.6734 0.0952 35.4345 0.0992 0.1716
giou box_count_recall boxwise lac multiplicative 14.4056 0.0262 25.0144 0.1000 35.4345 0.0992 0.1830
giou box_count_recall boxwise aps additive 14.4056 0.0262 28.6734 0.0952 34.8062 0.0953 0.1602
giou box_count_recall boxwise aps multiplicative 14.4056 0.0262 24.8396 0.0984 35.1776 0.0968 0.1805
hausdorff box_count_threshold thresholded lac additive 23.9744 0.0272 1.4441 0.0960 24.3973 0.1000 0.1728
hausdorff box_count_threshold thresholded lac multiplicative 23.9744 0.0272 1.0064 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold thresholded aps additive 23.9744 0.0272 1.4441 0.0924 23.4660 0.0984 0.1602
hausdorff box_count_threshold thresholded aps multiplicative 23.9744 0.0272 1.0064 0.0984 23.4660 0.0984 0.1797
hausdorff box_count_threshold pixelwise lac additive 23.9744 0.0272 1.0000 0.0984 24.3973 0.1000 0.1736
hausdorff box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold pixelwise aps additive 23.9744 0.0272 1.0000 0.0984 23.4660 0.0984 0.1602
hausdorff box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_threshold boxwise lac additive 23.9744 0.0272 1.3218 0.1048 24.3973 0.1000 0.1936
hausdorff box_count_threshold boxwise lac multiplicative 23.9744 0.0272 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold boxwise aps additive 23.9744 0.0272 1.3218 0.1048 23.4660 0.0984 0.1797
hausdorff box_count_threshold boxwise aps multiplicative 23.9744 0.0272 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_recall thresholded lac additive 14.4056 0.0262 1.4441 0.0960 24.3973 0.1000 0.1728
hausdorff box_count_recall thresholded lac multiplicative 14.4056 0.0262 1.0064 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall thresholded aps additive 14.4056 0.0262 1.4441 0.0924 23.4660 0.0984 0.1602
hausdorff box_count_recall thresholded aps multiplicative 14.4056 0.0262 1.0064 0.0984 23.4660 0.0984 0.1797
hausdorff box_count_recall pixelwise lac additive 14.4056 0.0262 1.0000 0.0984 24.3973 0.1000 0.1736
hausdorff box_count_recall pixelwise lac multiplicative 14.4056 0.0262 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall pixelwise aps additive 14.4056 0.0262 1.0000 0.0984 23.4660 0.0984 0.1602
hausdorff box_count_recall pixelwise aps multiplicative 14.4056 0.0262 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_recall boxwise lac additive 14.4056 0.0262 1.3218 0.1048 24.3973 0.1000 0.1936
hausdorff box_count_recall boxwise lac multiplicative 14.4056 0.0262 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall boxwise aps additive 14.4056 0.0262 1.3218 0.1048 23.4660 0.0984 0.1797
hausdorff box_count_recall boxwise aps multiplicative 14.4056 0.0262 1.0000 0.1016 23.4660 0.0984 0.1797
lac box_count_threshold thresholded lac additive 23.9744 0.0272 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_threshold thresholded lac multiplicative 23.9744 0.0272 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_threshold thresholded aps additive 23.9744 0.0272 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_threshold thresholded aps multiplicative 23.9744 0.0272 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_threshold pixelwise lac additive 23.9744 0.0272 10.9243 0.1000 0.6689 0.1008 0.1830
lac box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 6.6930 0.1016 0.6689 0.1008 0.1830
lac box_count_threshold pixelwise aps additive 23.9744 0.0272 10.9243 0.1000 0.8726 0.0780 0.1582
lac box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 6.6930 0.1016 0.8726 0.0780 0.1582
lac box_count_threshold boxwise lac additive 23.9744 0.0272 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_threshold boxwise lac multiplicative 23.9744 0.0272 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_threshold boxwise aps additive 23.9744 0.0272 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_threshold boxwise aps multiplicative 23.9744 0.0272 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_recall thresholded lac additive 14.4056 0.0262 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_recall thresholded lac multiplicative 14.4056 0.0262 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_recall thresholded aps additive 14.4056 0.0262 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_recall thresholded aps multiplicative 14.4056 0.0262 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_recall pixelwise lac additive 14.4056 0.0262 10.9243 0.1000 0.6689 0.1008 0.1830
lac box_count_recall pixelwise lac multiplicative 14.4056 0.0262 6.6930 0.1016 0.6689 0.1008 0.1830
lac box_count_recall pixelwise aps additive 14.4056 0.0262 10.9243 0.1000 0.8726 0.0780 0.1582
lac box_count_recall pixelwise aps multiplicative 14.4056 0.0262 6.6930 0.1016 0.8726 0.0780 0.1582
lac box_count_recall boxwise lac additive 14.4056 0.0262 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_recall boxwise lac multiplicative 14.4056 0.0262 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_recall boxwise aps additive 14.4056 0.0262 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_recall boxwise aps multiplicative 14.4056 0.0262 8.3136 0.1000 0.8726 0.0780 0.1582

### v2

@@ Appendix F Effect of Object Detection and Conformal Settings
mix box_count_threshold thresholded lac additive 22.6572 0.0268 2.5801 0.0976 0.9313 0.0878 0.1663
mix box_count_threshold thresholded lac multiplicative 22.6572 0.0268 1.7213 0.0884 0.9313 0.0878 0.1628
mix box_count_threshold thresholded aps additive 22.6572 0.0268 2.5801 0.0976 0.9936 0.0813 0.1609
mix box_count_threshold thresholded aps multiplicative 22.6572 0.0268 1.7213 0.0884 0.9936 0.0813 0.1582
mix box_count_threshold pixelwise lac additive 22.6572 0.0268 1.03 0.0981 0.9313 0.0878 0.1522
mix box_count_threshold pixelwise lac multiplicative 22.6572 0.0268 1.0179 0.0977 0.9313 0.0878 0.1524
mix box_count_threshold pixelwise aps additive 22.6572 0.0268 1.03 0.0981 0.9936 0.0813 0.149
mix box_count_threshold pixelwise aps multiplicative 22.6572 0.0268 1.0179 0.0977 0.9936 0.0813 0.1492
mix box_count_threshold boxwise lac additive 22.6572 0.0268 2.1558 0.0957 0.9313 0.0878 0.1477
mix box_count_threshold boxwise lac multiplicative 22.6572 0.0268 1.5515 0.0938 0.9313 0.0878 0.151
mix box_count_threshold boxwise aps additive 22.6572 0.0268 2.1558 0.0957 0.9936 0.0813 0.1439
mix box_count_threshold boxwise aps multiplicative 22.6572 0.0268 1.5515 0.0938 0.9936 0.0813 0.1477
mix box_count_recall thresholded lac additive 14.0464 0.0264 2.744 0.0988 1.064 0.0918 0.1675
mix box_count_recall thresholded lac multiplicative 14.0464 0.0264 1.9139 0.0768 1.064 0.0918 0.1538
mix box_count_recall thresholded aps additive 14.0464 0.0264 2.744 0.0988 1.0787 0.0918 0.1674
mix box_count_recall thresholded aps multiplicative 14.0464 0.0264 1.9139 0.0768 1.0787 0.0918 0.1537
mix box_count_recall pixelwise lac additive 14.0464 0.0264 1.0842 0.095 1.064 0.0918 0.1508
mix box_count_recall pixelwise lac multiplicative 14.0464 0.0264 1.0655 0.0967 1.064 0.0918 0.1538
mix box_count_recall pixelwise aps additive 14.0464 0.0264 1.0842 0.095 1.0787 0.0918 0.1508
mix box_count_recall pixelwise aps multiplicative 14.0464 0.0264 1.0655 0.0967 1.0787 0.0918 0.1538
mix box_count_recall boxwise lac additive 14.0464 0.0264 2.3195 0.0972 1.064 0.0918 0.1498
mix box_count_recall boxwise lac multiplicative 14.0464 0.0264 1.6957 0.0899 1.064 0.0918 0.1508
mix box_count_recall boxwise aps additive 14.0464 0.0264 2.3195 0.0972 1.0787 0.0918 0.1498
mix box_count_recall boxwise aps multiplicative 14.0464 0.0264 1.6957 0.0899 1.0787 0.0918 0.1509
hausdorff box_count_threshold thresholded lac additive 22.6572 0.0268 1.833 0.1024 25.3486 0.1012 0.19
hausdorff box_count_threshold thresholded lac multiplicative 22.6572 0.0268 1.3817 0.1112 25.3486 0.1012 0.2023
hausdorff box_count_threshold thresholded aps additive 22.6572 0.0268 1.833 0.1024 27.6101 0.1012 0.1901
hausdorff box_count_threshold thresholded aps multiplicative 22.6572 0.0268 1.3817 0.1112 27.6102 0.1012 0.2028
hausdorff box_count_threshold pixelwise lac additive 22.6572 0.0268 1.0009 0.0704 25.3486 0.1012 0.1536
hausdorff box_count_threshold pixelwise lac multiplicative 22.6572 0.0268 1.0179 0.0611 25.3486 0.1012 0.1463
hausdorff box_count_threshold pixelwise aps additive 22.6572 0.0268 1.0009 0.0704 27.6101 0.1012 0.1548
hausdorff box_count_threshold pixelwise aps multiplicative 22.6572 0.0268 1.0179 0.0611 27.6102 0.1012 0.1474
hausdorff box_count_threshold boxwise lac additive 22.6572 0.0268 1.6399 0.099 25.3486 0.1012 0.1742
hausdorff box_count_threshold boxwise lac multiplicative 22.6572 0.0268 1.2847 0.1014 25.3486 0.1012 0.1822
hausdorff box_count_threshold boxwise aps additive 22.6572 0.0268 1.6399 0.099 27.6101 0.1012 0.1752
hausdorff box_count_threshold boxwise aps multiplicative 22.6572 0.0268 1.2847 0.1014 27.6102 0.1012 0.1834
hausdorff box_count_recall thresholded lac additive 14.0464 0.0264 2.1505 0.1024 22.0349 0.0986 0.1871
hausdorff box_count_recall thresholded lac multiplicative 14.0464 0.0264 1.526 0.096 22.0349 0.0986 0.1852
hausdorff box_count_recall thresholded aps additive 14.0464 0.0264 2.1505 0.1024 23.6444 0.0989 0.1881
hausdorff box_count_recall thresholded aps multiplicative 14.0464 0.0264 1.526 0.096 23.6444 0.0989 0.1859
hausdorff box_count_recall pixelwise lac additive 14.0464 0.0264 0.999 0.0914 22.0349 0.0986 0.1677
hausdorff box_count_recall pixelwise lac multiplicative 14.0464 0.0264 1.017 0.0813 22.0349 0.0986 0.1597
hausdorff box_count_recall pixelwise aps additive 14.0464 0.0264 0.999 0.0914 23.6444 0.0989 0.169
hausdorff box_count_recall pixelwise aps multiplicative 14.0464 0.0264 1.017 0.0813 23.6444 0.0989 0.161
hausdorff box_count_recall boxwise lac additive 14.0464 0.0264 1.7532 0.1016 22.0349 0.0986 0.1739
hausdorff box_count_recall boxwise lac multiplicative 14.0464 0.0264 1.4049 0.1 22.0349 0.0986 0.1763
hausdorff box_count_recall boxwise aps additive 14.0464 0.0264 1.7532 0.1016 23.6444 0.0989 0.1748
hausdorff box_count_recall boxwise aps multiplicative 14.0464 0.0264 1.4049 0.1 23.6444 0.0989 0.1778
lac box_count_threshold thresholded lac additive 22.6572 0.0268 17.2748 0.0824 0.6533 0.0989 0.1694
lac box_count_threshold thresholded lac multiplicative 22.6572 0.0268 12.0794 0.1056 0.6533 0.0989 0.1872
lac box_count_threshold thresholded aps additive 22.6572 0.0268 17.2748 0.0824 0.9936 0.0548 0.1292
lac box_count_threshold thresholded aps multiplicative 22.6572 0.0268 12.0794 0.1056 0.9936 0.0548 0.1515
lac box_count_threshold pixelwise lac additive 22.6572 0.0268 12.277 0.0919 0.6533 0.0989 0.1594
lac box_count_threshold pixelwise lac multiplicative 22.6572 0.0268 7.7858 0.0999 0.6533 0.0989 0.1656
lac box_count_threshold pixelwise aps additive 22.6572 0.0268 12.277 0.0919 0.9936 0.0548 0.1268
lac box_count_threshold pixelwise aps multiplicative 22.6572 0.0268 7.7858 0.0999 0.9936 0.0548 0.1355
lac box_count_threshold boxwise lac additive 22.6572 0.0268 15.4127 0.0901 0.6533 0.0989 0.1565
lac box_count_threshold boxwise lac multiplicative 22.6572 0.0268 9.1685 0.0984 0.6533 0.0989 0.1631
lac box_count_threshold boxwise aps additive 22.6572 0.0268 15.4127 0.0901 0.9936 0.0548 0.1232
lac box_count_threshold boxwise aps multiplicative 22.6572 0.0268 9.1685 0.0984 0.9936 0.0548 0.1328
lac box_count_recall thresholded lac additive 14.0464 0.0264 15.1958 0.0796 0.817 0.1014 0.1689
lac box_count_recall thresholded lac multiplicative 14.0464 0.0264 12.9423 0.0956 0.817 0.1014 0.1805
lac box_count_recall thresholded aps additive 14.0464 0.0264 15.1958 0.0796 0.9928 0.0699 0.1399
lac box_count_recall thresholded aps multiplicative 14.0464 0.0264 12.9423 0.0956 0.9928 0.0699 0.1537
lac box_count_recall pixelwise lac additive 14.0464 0.0264 10.9039 0.0887 0.817 0.1014 0.1586
lac box_count_recall pixelwise lac multiplicative 14.0464 0.0264 8.2643 0.0943 0.817 0.1014 0.1641
lac box_count_recall pixelwise aps additive 14.0464 0.0264 10.9039 0.0887 0.9928 0.0699 0.1356
lac box_count_recall pixelwise aps multiplicative 14.0464 0.0264 8.2643 0.0943 0.9928 0.0699 0.1415
lac box_count_recall boxwise lac additive 14.0464 0.0264 13.6096 0.0852 0.817 0.1014 0.155
lac box_count_recall boxwise lac multiplicative 14.0464 0.0264 9.8398 0.0912 0.817 0.1014 0.1602
lac box_count_recall boxwise aps additive 14.0464 0.0264 13.6096 0.0852 0.9928 0.0699 0.1314
lac box_count_recall boxwise aps multiplicative 14.0464 0.0264 9.8398 0.0912 0.9928 0.0699 0.1376
giou box_count_threshold thresholded lac additive 22.6572 0.0268 32.7365 0.0524 34.0559 0.0876 0.1366
giou box_count_threshold thresholded lac multiplicative 22.6572 0.0268 147.1712 0.0468 34.0559 0.0876 0.1337
giou box_count_threshold thresholded aps additive 22.6572 0.0268 32.7365 0.0524 33.8632 0.0898 0.1389
giou box_count_threshold thresholded aps multiplicative 22.6572 0.0268 147.1712 0.0468 33.8633 0.0898 0.1359
giou box_count_threshold pixelwise lac additive 22.6572 0.0268 27.477 0.0681 34.0559 0.0876 0.1469
giou box_count_threshold pixelwise lac multiplicative 22.6572 0.0268 101.6635 0.0525 34.0559 0.0876 0.1374
giou box_count_threshold pixelwise aps additive 22.6572 0.0268 27.477 0.0681 33.8632 0.0898 0.1491
giou box_count_threshold pixelwise aps multiplicative 22.6572 0.0268 101.6635 0.0525 33.8633 0.0898 0.1394
giou box_count_threshold boxwise lac additive 22.6572 0.0268 31.6264 0.0573 34.0559 0.0876 0.136
giou box_count_threshold boxwise lac multiplicative 22.6572 0.0268 116.7034 0.05 34.0559 0.0876 0.1351
giou box_count_threshold boxwise aps additive 22.6572 0.0268 31.6264 0.0573 33.8632 0.0898 0.1384
giou box_count_threshold boxwise aps multiplicative 22.6572 0.0268 116.7034 0.05 33.8633 0.0898 0.1371
giou box_count_recall thresholded lac additive 14.0464 0.0264 28.1762 0.0312 32.523 0.0827 0.1122
giou box_count_recall thresholded lac multiplicative 14.0464 0.0264 153.597 0.0172 32.523 0.0827 0.0994
giou box_count_recall thresholded aps additive 14.0464 0.0264 28.1762 0.0312 32.335 0.0819 0.1113
giou box_count_recall thresholded aps multiplicative 14.0464 0.0264 153.597 0.0172 32.3353 0.0819 0.0986
giou box_count_recall pixelwise lac additive 14.0464 0.0264 23.6903 0.0541 32.523 0.0827 0.1297
giou box_count_recall pixelwise lac multiplicative 14.0464 0.0264 104.6357 0.0243 32.523 0.0827 0.1056
giou box_count_recall pixelwise aps additive 14.0464 0.0264 23.6903 0.0541 32.335 0.0819 0.1289
giou box_count_recall pixelwise aps multiplicative 14.0464 0.0264 104.6357 0.0243 32.3353 0.0819 0.1048
giou box_count_recall boxwise lac additive 14.0464 0.0264 27.2076 0.0418 32.523 0.0827 0.1187
giou box_count_recall boxwise lac multiplicative 14.0464 0.0264 120.4633 0.0215 32.523 0.0827 0.1029
giou box_count_recall boxwise aps additive 14.0464 0.0264 27.2076 0.0418 32.335 0.0819 0.1177
giou box_count_recall boxwise aps multiplicative 14.0464 0.0264 120.4633 0.0215 32.3353 0.0819 0.102